Suppose you are the manager of a soccer stadium where all the tickets always have to be sold at the same price. Two matches are scheduled to be played during the next fortnight, the first between Sundowns and Pirates and the second between two first Division sides.(USE ARC PRICE ELASTICITY)
(a) Market research indicates that you can sell 40 00 tickets for the Sundowns-Pirates clash at R10 each, or 30 000 tickets at R20 each, Which option would you choose ? What is the price Elasticity of the demand for tickets for this particular game ?
(b)Likewise, market research indicates that you can sell 15 000 tickets for the first Division fixture at R10 each, or 5 000 tickets at R20 each. Which option would you choose ?What is the price elasticity of the demand for the tickets for this game ?
1
Expert's answer
2019-03-29T10:47:49-0400

(a) Ed = (40000 - 30000)/(10 - 20)*(10 + 20)/(40000 + 30000) = -1000*30/70000 = -3/7, so the demand is inelastic.

(b) Ed = (15000 - 5000)/(10 - 20)*(10 + 20)/(15000 + 5000) = -1000*30/20000 = -1.5, so the demand is elastic.